OpenStudy (anonymous):

IMPORTANT: please answer fast. evaluate x^-2y^0/x^3y^-2 when x =2 and y=5 A)25/2 B)25/32 C)50

OpenStudy (johnweldon1993):

I believe there is something missing?

OpenStudy (anonymous):

fixed it oops

OpenStudy (johnweldon1993):

Ah lol, \[\large \frac{y^2}{x^5}\] \[\large \frac{(5)^2}{(2)^5}\] \[\large \frac{25}{32}\]

OpenStudy (anonymous):

Thankyou!

OpenStudy (johnweldon1993):

Of course :)